MATLAB Question: Write a script for the problem pictured. Please send the script, not just the printout. Make sure your script allows you to input a value for n

so when prompted, in the command window, user types n value. The n value should be the f_k if k ==1 .... Please also send print outs if you can.

BELOW IS A SAMPLE ANSWER THAT DID NOT PUT THE INPUT SCRIPT NEEDED FOR THIS QUESTION TO BE CORRECT. YOU CAN USE IT TO COME UP WITH A SCRIPT THAT DOES USE AN N INPUT (like n = ( 'input n ');)

P4.1.11 Write a script that inputs a positive integer n and then generates a length-10 row vector f according to the following formula: 1n ifk=1 k-1 fk-1/2 if 10 k1 and fk- is even. Your seript should plot the points (1.f10, fio) using the star marker

I have named the function as fk.

The following code has to be copied into a file named fk.m

function y = fk(k)
if (k == 1)
y = 1; % ACCORDING TO THE PROBLEM, THIS SHOULD PRINT n, NOT 1.
elseif(k > 1 && k <= 10)
previous = fk(k-1); % get fk(k-1), i.e. value of fk for k-1
if(mod(previous, 2) == 1) %if fk-1 is odd
y = 3 * previous + 1;
else %fk-1 is even
y = previous / 2;
end
else
y = 0;
end
end

============================

Now in a different file you copy this code

x = 1:10;
y = []; %start with an empty vector

%iterate over x value
for k = 1:10
y = [y , fk(k)] ; %append the value of function to vector
end

%now use x and y values and star marker to plot the graph
plot(x, y, '*')

The above code will produce the following output
Answer #1

The below code satisfies the given requirements as follows:

--------------------------------------------------------------------------

function y = fk(n) %%creating a function named fk which takes input as n and output a vector y

k=1 %%Initializing k value with 1

while (k >= 1 && k <= 10) %%while loop for k 10 times

if (k==1) %% if k is 1 y will be n

y(1)=n

else

if (mod(previous, 2) == 1) %%if the fk(k-1) is odd

y(k) = (3 * previous) + 1

elseif (mod(previous, 2) ~= 1) %%if the fk(k-1) is even

y(k) = previous / 2

else

y=0

end

end

k=k+1 %%increasing k value with 1

previous = y(k-1) %%storing the previous value into variable previous

end

end

For example, I am initializing the n value with 2 and plotting the required values to the graph with * as follows:

n=2

y = fk(n)

x = 1:10;

plot(x, y,'*')

xlabel("X-Axis")

ylabel("Y-Axis")

Screenshot of Output and code as in Matlab Edditor:-